OpenStudy (anonymous):

what is the product of 5 and x if x is 1

OpenStudy (anonymous):

What is the translation for "product of" ? does that mean multiplication or division ?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Well you have to translate this to find the answer , okay so product of means to factors being multiplied , so 5x , x=1 so 5times 1 = ??

OpenStudy (anonymous):

two*

OpenStudy (anonymous):

Product basically means multiplication. So,5*x=5*1=1+1+1+1+1=5 which is the correct answer as orangess mentioned.